OpenStudy (anonymous):

V = 100 sin (200πt + π/4) need to know the time for voltage to first reach a maximum

OpenStudy (helder_edwin):

can you use derivatives?

OpenStudy (phi):

what is the maximum value of sin(x)? what is x when sin(x) is at that maximum?

OpenStudy (anonymous):

tried this but think it is wrong dv/dt = (2 x 104 π) cos (200 π t + π/4) maximum occurs when dv/dt = 0 and d2v/dt2 is negative d2v/dt2 = -(4 x 106 π2) sin (200 π t + π/4) cos (200 π t + π/4) = 0 when t = 0 and when t = 0 d2v/dt2 is negative when t = 0, v = 100 as sin (200 π + π/t) = 1 dv/dt is a maximum when cos (200 π t + π/4) is either 1 or -1 which occurs when 200 π t + π/4 = 0 or π radians t is either = -1/800 or 1/800 seconds t = -0.00125 or 0.00125 seconds

OpenStudy (anonymous):

too much work max of sine is one when input is \(\frac{\pi}{2}\) set the inside part equal to \(\frac{\pi}{2}\) and solve for \(t\)

OpenStudy (helder_edwin):

\[ \large \frac{dV}{dt}=100\cos(200\pi t+\pi/4)200\pi \]

OpenStudy (asnaseer):

@tomtom - do you understand the reasoning used by @phi and @satellite73 ?

OpenStudy (anonymous):

not really

OpenStudy (asnaseer):

A sine curve looks like this: http://www.wolframalpha.com/input/?i=y%3Dsin%28x%29 notice its value always ranges between 1 and -1

OpenStudy (asnaseer):

so its maximum value is 1 - agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

now what would x need to be for sin(x) to be equal to 1?

OpenStudy (asnaseer):

are you aware of the fact that \(\sin(90^0)=1\)?

OpenStudy (asnaseer):

and that \(90^0=\frac{\pi}{2}\) radians

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so now all you need to do is to find what value of t will make this equal to 1:\[\sin(200\pi t+\frac{\pi}{4})=1\]which means we need to solve:\[200\pi t+\frac{\pi}{4}=\frac{\pi}{2}\]understand?

OpenStudy (anonymous):

yes -wow thank you

OpenStudy (asnaseer):

yw :)